4.2z=4(0.6z-1.2)
1
Expert's answer
2012-01-26T08:41:45-0500
opening the braces we get
4.2z=4*0.6z-4*1.2=2.4z-4.8

Substracting 2.4z from both parts of the equation we obtain
4.2z-2.4z=-4.8
1.8z=-4.8

Dividing both parts by 1.8 we obtain solution
z=-4.8/1.8=-2.(6)